# Signatures of Rapidly Rotating Stars with Chemically Homogeneous Evolution in the First Galaxies

### v1

@@ 3 Impact on First Galaxy UV Emission
Figure 2: Rest-frame UV luminosity per unit SFE $\eta_{\rm UV}$ for a top-hat star formation history within a timescale $t_{\rm SF}=100\ \rm Myr$. As in Fig. 1, the left and right panel show the results for Salpeter and top-heavy IMFs, with $\alpha=2.35$ and $1$, respectively. While the mass range of normal stars is fixed as $m_{\star}\in[1.7,300]\ \rm M_{\odot}$, we plot $\eta_{\rm UV}$ as a function of $m_{\star,\min}^{\rm CHE}$ for $f_{\rm CHE}\in[0,1]$, where $f_{\rm CHE}$ increases for lighter colors (from bottom to top). The dotted curve with errorbars shows the maximum $\eta_{\rm UV}$ in the alternative population model with a fixed overall IMF (corresponding to the maximum CHE mass fraction $f_{\rm CHE}^{\max}$), which is estimated to be in between $\eta_{\rm UV}(f_{\rm CHE}=f_{\rm CHE}^{\max})$ and $f_{\rm CHE}^{\max}\eta_{\rm UV}(f_{\rm CHE}=1)$. We also show the reference value $\eta_{\rm UV}\simeq 8.70\times 10^{27}$ from Madau & Dickinson (2014) with the dashed line, which is very close to our result for $f_{\rm CHE}=0$ given $\alpha=2.35$. The shaded regions show the results for Pop III stars with extremely top-heavy IMFs from Inayoshi et al. (2022).
@@
${\rm SFR/(M_{\odot}\ yr^{-1})}=\kappa_{\rm UV}L_{\nu}(\rm UV)/(erg\ s^{-1}\ Hz^{-1})\ .$ (1)
A reference value from local calibrations is $\eta_{\rm UV}\simeq 8.70\times 10^{27}$, corresponding to $\kappa_{\rm UV}=1.15\times 10^{-28}$ (Madau & Dickinson, 2014). Given the single-age starburst results (see Fig. 1), we calculate $\eta_{\rm UV}$ as a function of the CHE fraction $f_{\rm CHE}$ and minimum mass $m_{\star,\min}^{\rm CHE}$ under different IMFs by integrating the stellar mass-normalized light curve $\mathcal{L}_{\nu}(t)$ over a simple top-hat star formation history with a constant SFR for a period of $t_{\rm SF}$:
@@
We adopt $t_{\rm SF}=100$ Myr by default (see Appendix A for the results for $t_{\rm SF}=10$ and 1000 Myr) as the typical timescale of star formation probed by rest-frame UV emission, which is also close to the halo dynamical timescale and baryon cycle timescale at $z\sim 10$ (e.g., Anglés-Alcázar et al., 2017; Flores Velázquez et al., 2021; Tacchella et al., 2020). We have ignored the contribution of nebular emission and dust attenuation. It is shown in the Yggdrasil models (Zackrisson et al., 2011) that nebular emission in rest-frame UV is strongest for very young star-forming clouds, reaching up to $\sim 40\%$ of the stellar luminosity in moderately metal-enriched environments with $Z\gtrsim 0.0004\sim 0.03\ \rm Z_{\odot}$. However, it decays with time and becomes negligible ($\lesssim 10\%$ of the stellar contribution) shortly ($\sim$ a few Myr $\ll t_{\rm SF}$) after the starburst. Therefore, nebular emission is expected to be unimportant for the rest-frame UV emission of luminous high-$z$ galaxies observable by JWST [^4] Nebular emission can dominate the rest-frame UV luminosity (reaching a few times of the stellar contribution) in the first few Myr after a starburst and remain comparable to stellar emission for up to $\sim 20$ Myr for pristine (metal-free) star-forming clouds hosting massive stars, if leakage of Lyman-continuum photons is not significant (Zackrisson et al., 2011). We do not consider such scenarios, since it is unclear whether the nebula can survive under strong radiative feedback and there is no confident detection of galaxies dominated by metal-free stellar populations.. The effect of dust attenuation can be effectively absorbed into the UV variability parameter (see below and Shen et al., 2023).
@@
The results are shown in Fig. 2. We obtain $\eta_{\rm UV}\simeq 8.17\times 10^{27}$ in the canonical case of Salpeter IMF without CHE ($f_{\rm CHE}=0$), which is very close to the reference value $\eta_{\rm UV}\simeq 8.70\times 10^{27}$ (Madau & Dickinson, 2014). This further justifies our choice of $t_{\rm SF}=100$ Myr (as well as $m_{\star,\min}=1.7\ \rm M_{\odot}$). Interestingly, the enhancement of UV emission by CHE peaks at $m_{\star,\min}^{\rm CHE}=4\ (3)\ \rm M_{\odot}$ for the Salpeter (top-heavy) IMF, reaching $\eta_{\rm UV}\sim 5.14\ (2.73)\times 10^{28}$ for $f_{\rm CHE}=1$, which is higher than the reference value by a factor of $\sim 6$ (3). Here, $\eta_{\rm UV}$ does not increase monotonically with $m_{\star,\min}^{\rm CHE}$ because the (lifetime-integrated) rest-frame UV emission is only boosted by CHE for $m_{\star}\lesssim 100\ \rm M_{\odot}$ but reduced for more massive stars. The latter become too hot under CHE, producing primarily more energetic (ionizing) photons (Sibony et al., 2022). For the same reason, the effect of CHE is generally stronger at $m_{\star,\min}^{\rm CHE}\lesssim 10\ \rm M_{\odot}$ when $\alpha$ or $t_{\rm SF}$ increase (see Appendix A) or with a smaller IMF upper mass limit $m_{\star,\max}$. Without CHE, we obtain $\eta_{\rm UV}=1.42\times 10^{28}$ under our top-heavy IMF (that extends down to $m_{\star,\min}=1.7\ \rm M_{\odot}$ with $\alpha=1$), moderately increased by $\sim 74$ (63)compared with the Salpeter (reference) case. It is shown in Inayoshi et al. (2022, see their table. 1) that considering extremely top-heavy IMFs ($m_{\star}\gtrsim 10\ \rm M_{\odot}$) for Pop III stars further increases the UV emission efficiency to $\eta_{\rm UV}\sim 2.68-3.57\times 10^{28}$ (see also Bromm et al., 2001). Comparable and higher values of $\eta_{\rm UV}$ can be achieved by CHE with $f_{\rm CHE}\gtrsim 0.5$ and $m_{\star,\min}^{\rm CHE}\sim 2-10\ \rm M_{\odot}$ under a Salpeter IMF ($\alpha=2.35$).
@@
Given the results of $\eta_{\rm UV}(f_{\rm CHE})$ in the default model, the maximum UV efficiency in the alternative case of a fixed overall IMF should be in between $\eta_{\rm UV}(f_{\rm CHE}=f_{\rm CHE}^{\max})$ (with optimistic contributions from non-rotating stars below $m_{\star,\min}^{\rm CHE}$) and $f_{\rm CHE}^{\max}\eta_{\rm UV}(f_{\rm CHE}=1)$ (ignoring non-rotating stars below $m_{\star,\min}^{\rm CHE}$). The resulting values are very close to $\eta_{\rm UV}(f_{\rm CHE}=1)$ for the top-heavy IMF ($\alpha=1$) as $f_{\rm CHE}^{\max}>0.94$. For the Salpeter IMF with $f_{\rm CHE}^{\max}\sim 0.3-1$, the maximum UV emission efficiency peaks around $m_{\star,\min}^{\rm CHE}\sim 3-4\ \rm M_{\odot}$, reaching $\eta_{\rm UV}\sim 3.6-3.8\times 10^{28}$ given $f_{\rm CHE}^{\max}\sim 0.7-0.8$, which is still a factor of $\sim 4$ larger than the reference value.
@@
Once $\eta_{\rm UV}$ is known, we derive the UV luminosity function (UVLF) using the model in Shen et al. (2023). First, we calculate the median UV magnitude $M_{\rm UV}$ as a function of halo mass $M_{\rm h}$ from the median ${\rm SFR}=\epsilon_{\star}f_{\rm b}\dot{M}_{\rm h}$ estimated by the median halo accretion rate (Fakhouri et al., 2010)
@@
$\begin{split}\dot{M}_{\rm h}&\simeq 25.3\ {\rm M_{\odot}\ yr^{-1}}\left(\frac{M_{\rm h}}{10^{12}{\rm M_{\odot}}}\right)^{1.1}\\ &\times(1+1.65z)\sqrt{\Omega_{\rm m}(1+z)^{3}+\Omega_{\Lambda}}\ ,\end{split}$ (3)
@@
$\epsilon_{\star}=\frac{2\epsilon_{0}}{(M_{\rm h}/M_{0})^{-\gamma}+(M_{\rm h}/M_{0})^{\beta}}\ .$ (4)
@@
Here $f_{\rm b}\equiv\Omega_{\rm b}/\Omega_{\rm m}\simeq 0.16$ is the cosmic baryon fraction, and we adopt $\epsilon_{0}=0.1$, $\gamma=0.6$, and $\beta=0.5$ based on Behroozi et al. (2019, for the normalization and low-mass slope, matching the ${\rm SFR}$-$M_{\rm h}$ relation at $z=7$) and Harikane et al. (2022, for the high-mass slope). Next, this median $M_{\rm UV}$-$M_{\rm h}$ relation is combined with the halo mass function [^5] We use the hmf package (Murray et al., 2013) to calculate the halo mass function with the Tinker et al. (2010) model and real-space top-hat window function for a flat $\Lambda$CDM cosmology. The relevant cosmological parameters are based on Planck Collaboration et al. (2020) results: $H_{0}=67.66\ \rm km\ s^{-1}\ Mpc^{-1}$, $\Omega_{\rm m}=0.30966$, $n_{\rm s}=0.9667$, $\sigma_{8}=0.8159$, and $\Omega_{\rm b}=0.04897$. $dn/d\log M_{\rm h}$ to derive the median UVLF $dn_{\rm median}/dM_{\rm UV}=(dn/d\log M_{\rm h})|\frac{d\log M_{\rm h}}{dM_{\rm UV}}|$, which is then convolved with a Gaussian kernel of width $\sigma_{\rm UV}$ (mag) to predict the observable UVLF $\Phi\equiv dn/dM_{\rm UV}$. The last step captures the stochastic scatter in the scaling relations used above into the UV variability parameter $\sigma_{\rm UV}$.
@@
Figure 3: Constraints on UV emission efficiency and variability from JWST data. The colored curves show the values of $\epsilon_{0}\eta_{\rm UV}$ and $\sigma_{\rm UV}$ required to reproduce $\log(\Phi(M_{\rm UV}=-20.5)\ [\rm mag^{-1}\ cMpc^{-3}])=-5$ and $-5.3$ at $z=10$ (solid) and 12 (dash-dotted) based on spectroscopy, and $\log(\Phi(M_{\rm UV}=-20.5)\ [\rm mag^{-1}\ cMpc^{-3}])=-4.5$, $-5$, and $-5.2$ at $z=10$ (dashed), 12 (dotted), and 16 (long dashed) from photometry. Three representative cases of different $\eta_{\rm UV}$ with fixed $\epsilon_{0}=0.1$ are shown by the thick vertical lines: the top-heavy models with no CHE ($f_{\rm CHE}=0$, dashed) and optimal CHE ($f_{\rm CHE}=1$, $m_{\star,\min}^{\rm CHE}=3\ \rm M_{\odot}$, dash-dotted) and the Salpeter model with optimal CHE ($f_{\rm CHE}=1$, $m_{\star,\min}^{\rm CHE}=4\ \rm M_{\odot}$, dotted). The thin solid and dashed horizontal lines show the minimum and maximum UV variability estimated by Shen et al. (2023).
@@
Figure 4: UVLF for select degenerate models that meet the photometric constraints from JWST at $z=12$, i.e., $\log(\Phi(M_{\rm UV}=-20.5)\ [\rm mag^{-1}\ cMpc^{-3}])=-5$. These include the reference case with $\eta_{\rm UV}\simeq 8.70\times 10^{27}$ (solid) and the three representative models with enhanced UV emission from CHE and/or a top-heavy IMF considered in Fig. 3. For comparison, the observational data for $z\sim 11-13$ (Naidu et al., 2022; Bouwens et al., 2023; Harikane et al., 2023; Pérez-González et al., 2023; Adams et al., 2024; Casey et al., 2024; Donnan et al., 2023, 2024; Finkelstein et al., 2024; McLeod et al., 2024; Robertson et al., 2024; Willott et al., 2024) are shown using CoReCon (Garaldi, 2023), among which the data points with arrows are upper limits. We also plot the results for constant SFE $\epsilon_{\star}=0.03$, 0.1, 0.3, and 1 (from bottom to top) without UV variability as the thick solid lines.
@@
As illustrated in Shen et al. (2023, see their fig. 3), the UVLFs inferred from JWST observations place constraints on $\epsilon_{0}\eta_{\rm UV}$ and $\sigma_{\rm UV}$, which can be estimated by considering the value of $\Phi$ at a typical magnitude $M_{\rm UV}=-20.5$. Here, we require $\log(\Phi(M_{\rm UV}=-20.5)\ [\rm mag^{-1}\ cMpc^{-3}])=-5$ and $-5.3$ at $z=10$ and 12 according to spectroscopic observations, and $\log(\Phi(M_{\rm UV}=-20.5)\ [\rm mag^{-1}\ cMpc^{-3}])=-4.5$, $-5$, and $-5.2$ at $z=10$, 12, and 16 from photometric results, using the observational data compiled by the python package CoReCon [^6] https://corecon.readthedocs.io/en/latest/index.html (Garaldi, 2023). The corresponding values of $\epsilon_{0}\eta_{\rm UV}$ and $\sigma_{\rm UV}$ are shown in Fig. 3, where larger $\epsilon_{0}\eta_{\rm UV}$ and/or $\sigma_{\rm UV}$ are required to explain observations at higher $z$, and a smaller $\epsilon_{0}\eta_{\rm UV}$ can be compensated with a larger $\sigma_{\mathrm{UV}}$.
@@
To appreciate the trend and the degeneracy, we focus on 3 representative cases with boosted UV emission efficiency and fixed $\epsilon_{0}=0.1$: the top-heavy ($\alpha=1$) models with no CHE ($f_{\rm CHE}=0$) and optimal CHE ($f_{\rm CHE}=1$, $m_{\star,\min}^{\rm CHE}=4\ \rm M_{\odot}$), and the Salpeter ($\alpha=2.35$) model with optimal CHE ($f_{\rm CHE}=1$, $m_{\star,\min}^{\rm CHE}=3\ \rm M_{\odot}$). The corresponding UV variability parameters inferred from the photometric constraints at $z\sim 12$ (16) are $\sigma_{\rm UV}=1.75$, 1.40, and 0.98 (2.55, 2.31, and 2.06), respectively, while the reference value $\eta_{\rm UV}\simeq 8.70\times 10^{27}$ requires $\sigma_{\rm UV}=1.98$ (2.72). Clearly, with enhanced $\eta_{\rm UV}$ from CHE, smaller (less extreme) values of $\epsilon_{0}$ and $\sigma_{\rm UV}$ are required to explain JWST results. In particular, only for the Salpeter model with optimal CHE ($f_{\rm CHE}=1$, $m_{\star,\min}^{\rm CHE}=3\ \rm M_{\odot}$, $\alpha=2.35$) the $\sigma_{\rm UV}$ needed at $z\sim 16$ remains below the maximum value $2.2$ estimated by Shen et al. (2023). Fig. 4 shows the UVLFs of the (degenerate) models at $z\sim 12$, which all agree well with current observations, but differ slightly in shape: The evolution of $\Phi$ with $M_{\rm UV}$ is stronger for larger (smaller) $\epsilon_{0}\eta_{\rm UV}$ ($\sigma_{\rm UV}$), especially at the luminous end. Therefore, high precision measurements of UVLFs (in larger survey volumes) can potentially break the degeneracy between $\epsilon_{0}\eta_{\rm UV}$ and $\sigma_{\rm UV}$. Additional constraints from galaxy clustering properties are also useful (Muñoz et al., 2023; Sun et al., 2024).

### v2

@@ 3 Impact on First Galaxy UV Emission
${\rm SFR/(M_{\odot}~{}yr^{-1})}=\kappa_{\rm UV}L_{\nu}(\rm UV)/(erg~{}s^{-1}~{}Hz^{-1})~{}.$ (1)
A reference value from local calibrations adopted by Madau & Dickinson (2014) is $\eta_{\rm UV}\simeq 8.70\times 10^{27}$, corresponding to $\kappa_{\rm UV}=1.15\times~{}10^{-28}$. Given the single-age starburst results (see Fig. 1), we calculate $\eta_{\rm UV}$ as a function of the CHE fraction $f_{\rm CHE}$ and minimum mass $m_{\star,\min}^{\rm CHE}$ under different IMFs by integrating the stellar mass-normalized light curve $\mathcal{L}_{\nu}(t)$ over a simple top-hat star formation history with a constant SFR for a period of $t_{\rm SF}$:
@@
We adopt $t_{\rm SF}=100$ Myr by default (see Appendix A for the results for $t_{\rm SF}=10$ and 1000 Myr) as the typical timescale of star formation probed by rest-frame UV emission, which is also close to the halo dynamical timescale and baryon cycle timescale at $z\sim 10$ (e.g., Anglés-Alcázar et al., 2017; Flores Velázquez et al., 2021; Tacchella et al., 2020). We have ignored the contribution of nebular emission and dust attenuation. As discussed in Appendix B, for moderately metal-enriched ($Z\gtrsim 0.0004$) luminous high-$z$ galaxies observable by JWST, the enhancement of rest-frame UV emission by CHE with $t_{\rm SF}\gtrsim 10$ Myr can be underestimated by up to a factor of $\sim 2$ in the absence of nebular emission, which is also expected to be enhanced for CHE stars due to their strong ionizing power (Kubátová et al., 2019; Sibony et al., 2022), and the effect is stronger for more metal-poor systems (Zackrisson et al., 2011). Therefore, our results without nebular contributions should be regarded as conservative estimates. The effect of dust attenuation can be effectively absorbed into the UV variability parameter (see below and Shen et al., 2023).
@@
The results are shown in Fig. 2. We obtain $\eta_{\rm UV}\simeq~{}8.17~{}\times~{}10^{27}$ in the canonical case of Salpeter IMF without CHE ($f_{\rm CHE}=0$), which is very close to the reference value $\eta_{\rm UV}\simeq 8.70~{}\times~{}10^{27}$ (Madau & Dickinson, 2014). This further justifies our choice of $t_{\rm SF}=100$ Myr (as well as $m_{\star,\min}=1.7~{}\rm M_{\odot}$). Interestingly, the enhancement of UV emission by CHE peaks at $m_{\star,\min}^{\rm CHE}=4~{}(3)~{}\rm M_{\odot}$ for the Salpeter (top-heavy) IMF, reaching $\eta_{\rm UV}\sim 5.14~{}(2.73)~{}\times~{}10^{28}$ for $f_{\rm CHE}=1$, which is higher than the reference value by a factor of $\sim 6$ (3). Here, $\eta_{\rm UV}$ does not increase monotonically with $m_{\star,\min}^{\rm CHE}$ because the (lifetime-integrated) rest-frame UV emission of CHE stars has a peak around $m_{\star}\sim 2~{}\rm M_{\odot}$ and decreases for more massive stars, dropping even below the values for non-rotating stars at $m_{\star}\gtrsim 100~{}\rm M_{\odot}$. Here, massive stars become too hot under CHE, producing primarily more energetic (ionizing) photons (Sibony et al., 2022, Liu et al. in prep.). For the same reason, the effect of CHE is generally stronger at $m_{\star,\min}^{\rm CHE}\lesssim 10~{}\rm M_{\odot}$ when $\alpha$ or $t_{\rm SF}$ increase (see Appendix A) or with a smaller IMF upper mass limit $m_{\star,\max}$. Without CHE, we obtain $\eta_{\rm UV}=1.42~{}\times~{}10^{28}$ under our top-heavy IMF (that extends down to $m_{\star,\min}=1.7~{}\rm M_{\odot}$ with $\alpha=1$), moderately increased by $\sim 74$ (63)compared with the Salpeter (reference) case. It is shown in Inayoshi et al. (2022, see their table. 1) that considering extremely top-heavy IMFs ($m_{\star}\gtrsim 10~{}\rm M_{\odot}$) for Pop III stars further increases the UV emission efficiency to $\eta_{\rm UV}\sim 2.68-3.57~{}\times~{}10^{28}$ (see also Bromm et al., 2001). Comparable and higher values of $\eta_{\rm UV}$ can be achieved by CHE with $f_{\rm CHE}\gtrsim 0.5$ and $m_{\star,\min}^{\rm CHE}\sim 2-10~{}\rm M_{\odot}$ under a Salpeter IMF.
@@
Given the results of $\eta_{\rm UV}(f_{\rm CHE})$ in the default model, the maximum UV efficiency in the alternative case of a fixed overall IMF should be in between $\eta_{\rm UV}(f_{\rm CHE}=f_{\rm CHE}^{\max})$ (with optimistic contributions from non-rotating stars below $m_{\star,\min}^{\rm CHE}$) and $f_{\rm CHE}^{\max}\eta_{\rm UV}(f_{\rm CHE}=1)$ (ignoring non-rotating stars below $m_{\star,\min}^{\rm CHE}$). The resulting values are very close to $\eta_{\rm UV}(f_{\rm CHE}=1)$ for the top-heavy IMF ($\alpha=1$) as $f_{\rm CHE}^{\max}>0.94$. For the Salpeter IMF with $f_{\rm CHE}^{\max}\sim 0.3-1$, the maximum UV emission efficiency peaks around $m_{\star,\min}^{\rm CHE}\sim 3-4~{}\rm M_{\odot}$ given $f_{\rm CHE}^{\max}\sim 0.7-0.8$, reaching $\eta_{\rm UV}\sim 3.6-3.8~{}\times~{}10^{28}$, which is still a factor of $\sim 4$ larger than the reference value.
@@
Once $\eta_{\rm UV}$ is known, we derive the UV luminosity function (UVLF) using the model in Shen et al. (2023). First, we calculate the median UV magnitude $M_{\rm UV}$ as a function of halo mass $M_{\rm h}$ from the median ${\rm SFR}=~{}\epsilon_{\star}f_{\rm b}\dot{M}_{\rm h}$ estimated by the median halo accretion rate (Fakhouri et al., 2010)
@@
$\begin{split}\dot{M}_{\rm h}&\simeq 25.3~{}{\rm M_{\odot}~{}yr^{-1}}\left(\frac{M_{\rm h}}{10^{12}{\rm M_{\odot}}}\right)^{1.1}\\ &\times(1+1.65z)\sqrt{\Omega_{\rm m}(1+z)^{3}+\Omega_{\Lambda}}~{},\end{split}$ (3)
@@
$\epsilon_{\star}=\frac{2\epsilon_{0}}{(M_{\rm h}/M_{0})^{-\gamma}+(M_{\rm h}/M_{0})^{\xi}}~{}.$ (4)
@@
Here $f_{\rm b}\equiv\Omega_{\rm b}/\Omega_{\rm m}\simeq 0.16$ is the cosmic baryon fraction, and we adopt $\epsilon_{0}=0.1$, $\gamma=0.6$, and $\xi=0.5$ based on Behroozi et al. (2019, for the normalization and low-mass slope, matching the ${\rm SFR}$-$M_{\rm h}$ relation at $z=7$) and Harikane et al. (2022, for the high-mass slope). Next, this median $M_{\rm UV}$-$M_{\rm h}$ relation is combined with the halo mass function [^3] We use the hmf package (Murray et al., 2013) to calculate the halo mass function with the Tinker et al. (2010) model and real-space top-hat window function for a flat $\Lambda$CDM cosmology. The relevant cosmological parameters are based on Planck Collaboration et al. (2020) results: $H_{0}=67.66~{}\rm km~{}s^{-1}~{}Mpc^{-1}$, $\Omega_{\rm m}=0.30966$, $n_{\rm s}=0.9667$, $\sigma_{8}=0.8159$, and $\Omega_{\rm b}=0.04897$. $dn/d\log M_{\rm h}$ to derive the median UVLF $dn_{\rm median}/dM_{\rm UV}=(dn/d\log M_{\rm h})|\frac{d\log M_{\rm h}}{dM_{\rm UV}}|$, which is then convolved with a Gaussian kernel of width $\sigma_{\rm UV}$ (mag) to predict the observable UVLF $\Phi\equiv dn/dM_{\rm UV}$. The last step captures the stochastic scatter in the scaling relations used above into the UV variability parameter $\sigma_{\rm UV}$.
@@
Figure 3: Constraints on UV emission efficiency and variability from JWST data. The colored curves show the values of $\epsilon_{0}\eta_{\rm UV}$ and $\sigma_{\rm UV}$ required to reproduce $\log(\Phi(M_{\rm UV}=~{}-20.5)~{}[\rm mag^{-1}~{}cMpc^{-3}])=-5$ and $-5.3$ at $z=10$ (solid) and 12 (dash-dotted) based on spectroscopy, and $\log(\Phi(M_{\rm UV}=-20.5)~{}[\rm mag^{-1}~{}cMpc^{-3}])=-4.5$, $-5$, and $-5.2$ at $z=10$ (dashed), 12 (dotted), and 16 (long dashed) from photometry. Three representative cases of different $\eta_{\rm UV}$ with fixed $\epsilon_{0}=0.1$ are shown by the thick vertical lines: the top-heavy models with no CHE ($f_{\rm CHE}=0$, dashed) and optimal CHE ($f_{\rm CHE}=1$, $m_{\star,\min}^{\rm CHE}=3~{}\rm M_{\odot}$, dash-dotted) and the Salpeter model with optimal CHE ($f_{\rm CHE}=1$, $m_{\star,\min}^{\rm CHE}=4~{}\rm M_{\odot}$, dotted). The thin solid and dashed horizontal lines show the minimum and maximum UV variability estimated by Shen et al. (2023).
@@
Figure 4: UVLF for select degenerate models that meet the photometric constraints from JWST at $z=12$, i.e., $\log(\Phi(M_{\rm UV}=-20.5)~{}[\rm mag^{-1}~{}cMpc^{-3}])=-5$. These include the reference case with $\eta_{\rm UV}\simeq 8.70~{}\times~{}10^{27}$ (solid) and the three representative models with enhanced UV emission from CHE and/or a top-heavy IMF considered in Fig. 3. For comparison, the observational data for $z\sim 11-13$ (Naidu et al., 2022; Bouwens et al., 2023; Harikane et al., 2023; Pérez-González et al., 2023; Adams et al., 2024; Casey et al., 2024; Donnan et al., 2023, 2024; Finkelstein et al., 2024; McLeod et al., 2024; Robertson et al., 2024; Willott et al., 2024) are shown using CoReCon (Garaldi, 2023), among which the data points with arrows are upper limits. We also plot the results for constant SFE $\epsilon_{\star}=0.03$, 0.1, 0.3, and 1 (from bottom to top) without UV variability as the thick solid lines.
@@
As illustrated in Shen et al. (2023, see their fig. 3), the UVLFs inferred from JWST observations place constraints on $\epsilon_{0}\eta_{\rm UV}$ and $\sigma_{\rm UV}$, which can be estimated by considering the value of $\Phi$ at a typical magnitude $M_{\rm UV}=-20.5$. Here, we require $\log(\Phi(M_{\rm UV}=~{}-20.5)~{}[\rm mag^{-1}~{}cMpc^{-3}])=-5$ and $-5.3$ at $z=10$ and 12 according to spectroscopic observations, and $\log(\Phi(M_{\rm UV}=-20.5)\rm~{}[mag^{-1}~{}cMpc^{-3}])=-4.5$, $-5$, and $-5.2$ at $z=10$, 12, and 16 from photometric results, using the observational data compiled by the python package CoReCon [^4] https://corecon.readthedocs.io/en/latest/index.html (Garaldi, 2023). The corresponding values of $\epsilon_{0}\eta_{\rm UV}$ and $\sigma_{\rm UV}$ are shown in Fig. 3, where larger $\epsilon_{0}\eta_{\rm UV}$ and/or $\sigma_{\rm UV}$ are required to explain observations at higher $z$, and a smaller $\epsilon_{0}\eta_{\rm UV}$ can be compensated with a larger $\sigma_{\mathrm{UV}}$.
@@
To appreciate the trend and the degeneracy, we focus on 3 representative cases with boosted UV emission efficiency and fixed $\epsilon_{0}=0.1$: the top-heavy ($\alpha=1$) models with no CHE ($f_{\rm CHE}=0$) and optimal CHE ($f_{\rm CHE}=1$, $m_{\star,\min}^{\rm CHE}=4~{}\rm M_{\odot}$), and the Salpeter ($\alpha=2.35$) model with optimal CHE ($f_{\rm CHE}=1$, $m_{\star,\min}^{\rm CHE}=3~{}\rm M_{\odot}$). The corresponding UV variability parameters inferred from the photometric constraints at $z\sim 12$ (16) are $\sigma_{\rm UV}=1.75$, 1.40, and 0.98 (2.55, 2.31, and 2.06), respectively, while the reference value $\eta_{\rm UV}\simeq 8.70~{}\times~{}10^{27}$ requires $\sigma_{\rm UV}=1.98$ (2.72). Clearly, with enhanced $\eta_{\rm UV}$ from CHE, smaller (less extreme) values of $\epsilon_{0}$ and $\sigma_{\rm UV}$ are required to explain JWST results. In particular, only for the Salpeter model with optimal CHE ($f_{\rm CHE}=1$, $m_{\star,\min}^{\rm CHE}=3~{}\rm M_{\odot}$, $\alpha=2.35$) the $\sigma_{\rm UV}$ needed at $z\sim 16$ remains below the maximum value $2.2$ estimated by Shen et al. (2023). Fig. 4 shows the UVLFs of the (degenerate) models at $z\sim 12$, which all agree well with current observations, but differ slightly in shape: The evolution of $\Phi$ with $M_{\rm UV}$ is stronger for larger (smaller) $\epsilon_{0}\eta_{\rm UV}$ ($\sigma_{\rm UV}$), especially at the luminous end. Therefore, high precision measurements of UVLFs (in larger survey volumes) can potentially break the degeneracy between $\epsilon_{0}\eta_{\rm UV}$ and $\sigma_{\rm UV}$. Additional constraints from galaxy clustering properties are also useful (Muñoz et al., 2023; Sun et al., 2024).